• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Adaq

    Log entry time 13:41:11 on January17,2001

    Entry number 54905

    keyword=Target IOC reboot

    13:15 Graphs on target went flat, IOC rebooted. High power heat was not restored
    automatically. This was because the setpoint of the high power heater was 6.3K instead of 22K. Dave Meekins saved the day. R. Holt
    Also, we were getting high temperature alarms on the ESR. J.P. Chen set the
    High setpoint to 14.5 rather than 14.0
    13:30 Target is stable again.